OpenStudy (anonymous):

Solve this equation for x: log10^(x-1)=5

sam (.sam.):

x-1=10^5 x=10^5 +1 x=100001

sam (.sam.):

Using rule \[\huge \log_{a}b=c \leftarrow \rightarrow b=a^c\]